rods»view/4

Status Summary for ROD ID 4

<<Prev | Next>>

Test NameStatusNotesModified
Board Status 0 one bad connection between V2 and V4 bottom (channel 15) 12 June 2007: checked all other connections; all are GOOD except noted above 13 June: Brought to CERN shop; they reheated the V2 (with much skepticism) V2-> V4 connections now all good Noticed after closer examination of L1iD that 3 drivers VME to backplane connection are *missing*!?! Drivers replaced and board works 17 July 2007: board failed in pit (Mike saw it fail; go from working to half working to not working.) Channel 15 open from V2->V4 bottom again; ALSO test fifo line 58 is stuck high, which seems to cause test data not to work on channel 15 top as well. Real data does work on top fully unmasked; on bottom with usual masks 0x8000 on masks 0 and 2. Pushing on the V2 removes all these problems. 27 May 2010: Line VChipPins<42> low in the V4T, this is LB_A(8). Subsequently the V4T cannot be programmed correctly. Line V2Pins<75> low in the V4T, this might be an slink channel. Line V2Pins<15> low in V4B as mentioned above (Alex). 27-05-2010
fuses 1
12-12-2007
U31
12-12-2007
U32
12-12-2007
U33
12-12-2007
U34
12-12-2007
U35
12-12-2007
U81
12-12-2007
U82
12-12-2007
I5V_initial
12-12-2007
I3V initial
12-12-2007
proms load
12-12-2007
I5V proms
12-12-2007
I3V proms
12-12-2007
V4T temp
12-12-2007
V4B temp
12-12-2007
connectivity
12-12-2007
loop data
12-12-2007
slink
12-12-2007
dram
12-12-2007
nse communication
12-12-2007
sram
12-12-2007
optical loop
12-12-2007
patch panel
12-12-2007
qpll lock
12-12-2007
ttc connectivity
12-12-2007
test fifo
12-12-2007
ddr reset
12-12-2007
assembly 0
see history.
12-12-2007